# Top-down machine learning of coarse-grained protein force-fields

### v1

@@ 1 Introduction
In recent years, the adoption of machine learning algorithms in molecular dynamics has gained traction, driven by the increasing availability of experimental data, computational power, and auto-differentiation software20, 21, 22, 23. A particularly promising application of machine learning in MD is the development of neural network potentials (NNPs)24. NNPs can effectively model many-body interactions and represent potential energy surfaces25, and have been employed to construct machine-learned coarse-grained force fields for biomolecules26, 27, 22, 28, 29, 30.
The development of CG machine-learned NNPs for proteins generally adopts a bottom-up approach, which attempts to reproduce the reference fine-grained statistics. Initial studies focused on directly learning potentials from high-fidelity simulations through variational force-matching27, 22, 29, while more recent research has explored data-efficient strategies, such as flow-matching31 and denoising diffusion probabilistic models32. This bottom-up approach has the advantage of preserving the thermodynamics of the projected degrees of freedom. However, it also necessitates a large quantity of all-atom 3D conformations and their corresponding energies and forces sampled from the equilibrium distribution for training the machine learning model. This requirement can be computationally expensive and may result in poor extrapolation in regions of conformational and sequence space where data is scarce33.
On the other hand, an alternative approach for learning potential energy functions has been demonstrated by Greener et al.34 through the application of end-to-end differentiable molecular simulations (DMS). However, this method faces challenges when applied to medium-to-large proteins, as it requires a significant amount of memory to store all simulation operations during the forward pass, which are then used in the backward pass. This can lead to exploding gradients, causing instability during training as the accumulated gradients result in large updates in the neural network weights. To address this issue, the differentiable trajectory reweighting (DiffTre) method has been developed and applied to learn NNPs for atomistic systems in a more memory-efficient manner35. Bypassing the differentiation of the entire trajectory through the MD simulation for time-independent observables, DiffTre enables the learning of top-down coarse-grained potentials.
@@ 2.1 Coarse Grained Molecular Dynamics
The design of a coarse-grained model starts with the definition of the variables that should be preserved after the dimensionality reduction. In this study, to reduce the dimensionality of system $\textbf{R}\in\mathbb{R}^{3N}$, we establish a linear mapping $\Xi:\mathbb{R}^{3N}\rightarrow\mathbb{R}^{3n}$, projecting it onto a lower dimensional representation $r\in\mathbb{R}^{3n}$. We transform the all-atom representation into a $C_{\alpha}$ atom representation, with retained atoms referred to as coarse-grained ”beads”. Each $C_{\alpha}$ bead is assigned a bead type based on amino acid type, resulting in 21 unique bead types, which are identified by distinct integers.
Molecular dynamics simulations are employed for training and testing the NNP. We utilize TorchMD 28, a fully differentiable MD package written in PyTorch 36, enabling the execution of multiple parallel trajectories. To confine the space explored by the dynamics and incorporate physical information, we apply a prior energy function $U_{\lambda}(r,\phi)$. TorchMD integrates prior energy terms and the NNP to compute total potential energy. Consequently, the potential energy function is decomposed into a prior potential $U_{\lambda}$ and a neural network potential $U_{\theta}$, with the total potential energy given by:
@@
where $U_{\theta}$ represents the potential energy derived from the NNP, parameterized by parameters $\theta$. The network is a graph neural network with a SchNet-based 37 architecture, a continuous-filter convolution neural network capable of modeling atomic interactions. The implementation, available in the Torchmd-NET package 30, is defined by Majewski et al. 29. The prior energy $U_{\lambda}$ is parameterized by constant parameters $\lambda$, which can be decomposed into three terms: a pairwise bonded term to prevent chain breaking, a nonbonded repulsive term to avoid structure collapse into overlapping beads, and a dihedral term to enforce chirality in the system. The Supporting Information provides a detailed description of the prior energy terms. The total energy of the system is then expressed as:
@@ 2.2 Differentiable Trajectory Reweighting
We implement a version of DiffTre 35 in PyTorch to train the NNP, facilitating parallel and distributed training across multiple GPUs and nodes. The package is modular, allowing training for any molecular system and using any experimental observable as a training objective. DiffTre is used to match K outputs of molecular dynamics simulations to experimental observables. In this study, we focus on the folding of coarse-grained proteins, using conformations $\textbf{r}_{0}\in\mathbb{R}^{3n}$ of proteins in their native states as experimental observables, where $n$ denotes the number of beads in the system.
@@
As illustrated in Algorithm 1, we simulate multiple trajectories of length $N$ in parallel, sampling $K$ uncorrelated states $S_{i}$ from each trajectory $\{\textbf{S}_{i}\}_{i=1}^{N}$. For each state, we compute the root-mean square deviation (RMSD) between the state’s coordinates and the native conformation coordinates. Subsequently, the weighted RMSD ensemble average can be calculated as
$\langle RMSD\rangle=\sum_{i=1}^{N}w_{i}RMSD(\textbf{r}_{0},\textbf{r}_{i}),$ (3)
@@
where $U_{\lambda,\hat{\theta}}$ denotes the potential energy calculated with the reference parameters that generate the trajectory, $U_{\lambda,\theta}$ represents the potential energy calculated with the parameters to be updated, and $\beta=1/(k_{B}T)$, with $k_{B}$ being the Boltzmann constant and T the temperature. Note that before the first backward pass, $\theta=\hat{\theta}$, thus $w_{i}=1/N$.
@@ 2 $\hat{\theta}\longleftarrow\theta$
4      Draw a minibatch of samples { ${r_{0}^{1},\dots,r_{0}^{N}}$ from $p_{data}\sim\mathcal{D}$.
5      Sample $K$ states for every sample of the minibatch $\{S_{1},\dots,S_{K}\}_{1},\dots,\{S_{1},\dots,S_{K}\}_{N}$} from $q_{\lambda,\hat{\theta}}\propto exp(-U_{\lambda,\hat{\theta}})$ (Using Molecular Dynamics trajectories of length $N$)
@@ 6      Update the neural network parameters by stochastic gradient descent:
$\langle RMSD\rangle\simeq\sum_{i=1}^{N}w_{i}RMSD(\textbf{r}_{0},\textbf{r}_{i})$
$\nabla_{\theta}\mathcal{L}_{\theta}=\frac{1}{M}\sum^{M}_{k=1}[max(0,\langle RMSD\rangle+m)]$
@@ 2.3 Markov State Models
Markov State Models (MSM) 38, 39, 40, 41, 42 are employed to analyze the CG simulations and compare them with their corresponding all-atom simulations. MSMs can describe the entire dynamics of a system by partitioning it into $n$ discrete states. For a system to be Markovian, it must be ”memoryless”, meaning that future states depend only on the current state. In the case of Markovian systems, such as MD simulations, a transition probability matrix can be constructed, characterized by the $n$ states and the lag time $\tau$ at which the system’s state is recorded. From this matrix, state populations and conditional pairwise transition probabilities can be obtained, and the state populations can be converted into free energies.
In this work, we employ time-lagged independent component analysis (TICA) 43, 44 to project the high-dimensional conformational space into an optimally reduced low-dimensional space. Following this, the resulting space is discretized using K-means clustering for MSM construction. We featurize the simulation data into pairwise $C_{\alpha}$ distances and use TICA to project the data onto the first four components. For the coarse-grained (CG) data, we adopt the approach presented by Maciej et al. 29, using the covariance matrices of the all-atom molecular dynamics (MD) to project the first three components. This method ensures consistency with established methodologies and facilitates further analysis.
We use MSMs in the coarse-grained trajectories because we had them already for the all-atom trajectories and to evaluate the shape of the folding basins. However, we have no expectation that the thermodynamics or kinetics of these coarse-grained simulations have anything to do with the original ones as the training methods do not preserve these quantities. It can be interpreted as a way to get stable states, which we can take as predictions for the native structure. By comparing the most probable macrostate to the protein’s native conformation, we can evaluate the predictive capabilities of our CG model. For clustering the data we apply the Pairwise Constrained Component Analysis 45 (PCCA) algorithm.
@@ 2.4 Datasets
The first dataset comprises the crystal structures of 12 fast-folding proteins, previously studied by Lindorff-Larsen et al.46 (using all-atom MD) and Majewski et al 29 (using Machine Learned CG MD). These proteins exhibit a variety of secondary structure elements, including $\alpha$-helices and $\beta$-sheets.
The second dataset, used for training the general model, was created by searching the AlphaFold Database 47. This dataset contains 15,020 proteins between 50 and 150 residues, with computationally predicted structures solved using AlphaFold2 48. The selected structures in this dataset are high-quality predictions, with a global predicted local-distance difference test (pLDDT) higher than 90, and are clustered at 50% sequence identity using Usearch 49. We also removed any hits with more than 90% sequence identity to the fast-folding proteins used for testing. This approach ensures a diverse representation of protein structures in the dataset, allowing the NNP to generalize effectively to new and previously unseen protein structures. By incorporating AlphaFold2-predicted structures into our dataset, we significantly increase its size compared to using experimentally solved structures alone. The datasets are available in the supplementary information.
@@ 3 Results
We trained the FF-NNP using the dataset of 12 fast-folding proteins. For training, the learning rate was set to $\epsilon=0.0001$, and the loss function was defined by Equation (5). The simulation temperature was set to 298 K, and the timestep was 5 fs. We employed trajectories of 1024 steps, with 128 states used for reweighting each trajectory. The margin was set to $m=-1.0;\mathrm{\AA}$, and the mini-batch size was 12.
Compared to bottom-up approaches 29, 31, 32, 22, 27, our training process did not require generating expensive reference all-atom data beforehand, and the training took only 5 hours on a single NVIDIA GeForce RTX 2080 GPU. Despite this, the model successfully folded most of the proteins and stabilized their native conformations.
@@
FF-NNP G-NNP
Protein Min RMSD (Å) Mean RMSD (Å) Macro prob. (%) Min RMSD (Å) Mean RMSD (Å) Macro prob. (%)
Chignolin 0.3 0.9 $\pm$ 0.7 42.5 $\pm$ 0.2 0.2 1.5 $\pm$ 0.4 16.3 $\pm$ 0.2
Trp-cage 1.1 3.2 $\pm$ 1.4 5.5 $\pm$ 0.1 4.0 5.5 $\pm$ 0.5 11.0 $\pm$ 0.1
BBA 0.4 1.3 $\pm$ 0.7 30.0 $\pm$ 0.1 1.6 2.5 $\pm$ 0.4 42.9 $\pm$ 0.1
WW-domain 0.4 1.0 $\pm$ 0.6 1.8 $\pm$ 0.2 2.7 4.9 $\pm$ 0.9 4.4 $\pm$ 0.1
Villin 0.4 1.0 $\pm$ 0.5 18.6 $\pm$ 0.1 2.2 6.9 $\pm$ 1.4 6.9 $\pm$ 0.1
NTL9 0.5 0.9 $\pm$ 0.3 16.0 $\pm$ 0.1 2.7 4.7 $\pm$ 0.4 2.9 $\pm$ 0.1
BBL 0.5 1.8 $\pm$ 0.6 4.6 $\pm$ 0.2 3.3 6.1 $\pm$ 1.3 17.0 $\pm$ 0.1
Protein B 4.0 8.6 $\pm$ 2.5 9.9 $\pm$ 0.1 4.5 6.6 $\pm$ 0.7 4.8 $\pm$ 0.1
Homeodomain 0.5 3.6 $\pm$ 3.9 35.0 $\pm$ 0.1 4.1 7.2 $\pm$ 0.9 7.1 $\pm$ 0.1
Protein G 0.5 1.0 $\pm$ 0.4 1.5 $\pm$ 0.1 1.7 2.8 $\pm$ 0.4 6.1 $\pm$ 0.1
a3d 3.4 8.5 $\pm$ 2.3 5.8 $\pm$ 0.1 4.1 7.0 $\pm$ 2.3 5.7 $\pm$ 0.1
$\lambda$-repressor 4.9 6.8 $\pm$ 1.5 0.4 $\pm$ 0.2 4.6 7.5 $\pm$ 1.1 4.9 $\pm$ 0.1
Table 1: Minimum average RMSD macrostate statistic from MSM built with CG simulations of the fast-folders and general NNP. The data shows the average and minimum RMSD of the macrostate, as well as its equilibrium probabilities in percentage (Macro prob.) with standard deviation. In bold the proteins with Mean RMSD $<$ 3.0 Å.
@@
In the previous section, we demonstrated that a folding NNP can be learned for a small set of proteins. In this section, we present the results of the G-NNP, which is trained on a much larger dataset of 15 thousand protein structures. This enables us to test the generalization capabilities of our approach. We used a random 90/10% training/validation split. Furthermore, we built the training dataset such that it did not contain sequences with a sequence similarity greater than 90% to the fast folders. The closest homologs to fast folders present in the dataset after filtering are reported in Table (S2) in the Supporting Information. Thus, we use the 12 fast-folding proteins to provide a well-known but independent test set of our model’s ability to generalize to proteins that fall outside of the training set and allow us to directly compare the results with the FF-NNP.
For training the G-NNP, the learning rate was set to $\epsilon=0.0005$. We used trajectories of 100 steps, the number of states used to reweight each trajectory was 20, and the mini-batch size was set to 32. As for the big dataset we have much more data, we can set the margin to $m=0\mathrm{\AA}$ without suffering from overfitting. We found that this combination was more optimal for training on a large dataset.
@@
Results from the MSM analysis reveal that for specific proteins, such as Chignolin, BBA, and Protein G, the minimum RMSD macrostate aligns with the native structure, exhibiting an average RMSD of less than 3.0 Å. Moreover, their equilibrium probabilities are 16.3 $\pm$ 0.2 % for Chignolin, 42.9 $\pm$ 0.1 % for BBA, and 6.1 $\pm$ 0.1 % for Protein G (Table 1). This evidence suggests that our NNP can effectively generalize when trained on an ample dataset. For the other proteins, the minimum RMSD of the macrostates is consistently lower than 5 Å. As a result, although the macrostates do not precisely match the native structure, near-native conformations are sampled during the simulations with notable probability.
Simulations for Chignolin, BBA, and Homeodomain proteins successfully sampled folding events, even when starting from completely unfolded conformations. This indicates that the G-NNP can capture the underlying folding mechanisms that guide these proteins to adopt their native structures. The folding events observed in our simulations, illustrated in Figure 2, provide compelling evidence that our G-NNP generalizes beyond the training set, accurately modeling the folding behavior of proteins not included in the training set.

### v4

@@ 1 Introduction
Early efforts in the development of coarse-grained (CG) protein potentials laid the foundation for knowledge-based (KB) protein models. Early KB works 20, 21 established significant progress in creating statistical potentials from protein data bank (PDB) structures. Subsequent KB studies 22, 23, 24, 25, 26 furthered this field by designing CG potentials to stabilize PDB structures. Alternatively, bottom-up approaches for parameterizing a CG model from PDB conformations have also been utilized. These methodologies postulate that PDB conformations dominate the equilibrium ensemble so can be used to determine transferable interaction potentials for CG protein models, with statistical physics approaches that treat many-mody structural correlation 27, maximizing the likelihood that the CG model samples PDB conformations, as demonstrated in the Bayesian approach28 or with the relative entropy approach 29.
In recent years, the adoption of machine learning algorithms in molecular dynamics has gained traction, driven by the increasing availability of experimental data, computational power, and auto-differentiation software30, 31, 32, 33. A particularly promising application of machine learning in MD is the development of neural network potentials (NNPs)34. NNPs can effectively model many-body interactions and represent potential energy surfaces35, and have been employed to construct machine-learned coarse-grained force fields for biomolecules36, 37, 32, 38, 39, 40.
The development of CG machine-learned NNPs for proteins generally adopts a bottom-up approach, which attempts to reproduce the reference fine-grained statistics. Initial studies focused on directly learning potentials from high-fidelity simulations through variational force-matching37, 32, 39, while more recent research has explored data-efficient strategies, such as flow-matching41 and denoising diffusion probabilistic models42. This bottom-up approach has the advantage of preserving the thermodynamics of the projected degrees of freedom. However, it also necessitates a large quantity of all-atom 3D conformations and their corresponding energies and forces sampled from the equilibrium distribution for training the machine learning model. This requirement can be computationally expensive and may result in poor extrapolation in regions of conformational and sequence space where data is scarce43.
On the other hand, an alternative approach for learning potential energy functions has been demonstrated by Greener et al.44 through the application of end-to-end differentiable molecular simulations (DMS). However, this method faces challenges when applied to medium-to-large proteins, as it requires a significant amount of memory to store all simulation operations during the forward pass, which are then used in the backward pass. This can lead to exploding gradients, causing instability during training as the accumulated gradients result in large updates in the neural network weights. To address this issue, the differentiable trajectory reweighting (DiffTre) method has been developed and applied to learn NNPs for atomistic systems in a more memory-efficient manner45. Bypassing the differentiation of the entire trajectory through the MD simulation for time-independent observables, DiffTre enables the learning of top-down coarse-grained potentials.
@@ 2.1 Coarse Grained Molecular Dynamics
[…]
Molecular dynamics simulations are employed for training and testing the NNP. We utilize TorchMD 38, a fully differentiable MD package written in PyTorch 46, enabling the execution of multiple parallel trajectories. To confine the space explored by the dynamics and incorporate physical information, we apply a prior energy function $U_{\lambda}(r,\phi)$. TorchMD integrates prior energy terms and the NNP to compute total potential energy. Consequently, the potential energy function is decomposed into a prior potential $U_{\lambda}$ and a neural network potential $U_{\theta}$, with the total potential energy given by:
@@
where $U_{\theta}$ represents the potential energy derived from the NNP, parameterized by parameters $\theta$. The network is a graph neural network with a SchNet-based 47 architecture, a continuous-filter convolution neural network capable of modeling atomic interactions. The implementation, available in the Torchmd-NET package 40, is defined by Majewski et al. 39. The prior energy $U_{\lambda}$ is parameterized by constant parameters $\lambda$, which can be decomposed into three terms: a pairwise bonded term to prevent chain breaking, a nonbonded repulsive term to avoid structure collapse into overlapping beads, and a dihedral term to enforce chirality in the system. The Supporting Information provides a detailed description of the prior energy terms. The total energy of the system is then expressed as:
@@ 2.2 Differentiable Trajectory Reweighting
We implement a version of DiffTre 45 in PyTorch to train the NNP, facilitating parallel and distributed training across multiple GPUs and nodes. The package is modular, allowing training for any molecular system and using any experimental observable as a training objective. DiffTre is used to match K outputs of molecular dynamics simulations to experimental observables. In this study, we focus on the folding of coarse-grained proteins, using conformations $\textbf{r}_{0}\in\mathbb{R}^{3n}$ of proteins in their native states as experimental observables, where $n$ denotes the number of beads in the system.
@@
As illustrated in Algorithm 1, we simulate multiple trajectories of length $N$ in parallel, sampling $K$ uncorrelated states $S_{i}$ from each trajectory $\{S_{i}\}_{i=1}^{N}$. For each state, we compute the root-mean square deviation (RMSD) between the state’s coordinates and the native conformation coordinates. Subsequently, the weighted RMSD ensemble average can be calculated as
$\langle RMSD\rangle=\sum_{i=1}^{K}w_{i}RMSD(\textbf{r}_{0},\textbf{r}_{i}),$ (3)
@@
where $U_{\lambda,\hat{\theta}}$ denotes the potential energy calculated with the reference parameters that generate the trajectory, $U_{\lambda,\theta}$ represents the potential energy calculated with the parameters to be updated, and $\beta=1/(k_{B}T)$, with $k_{B}$ being the Boltzmann constant and T the temperature. Note that before the first backward pass, $\theta=\hat{\theta}$, thus $w_{i}=1/K$.
@@ 2 $\hat{\theta}\longleftarrow\theta$
4      Draw a minibatch of samples { ${r_{0}^{1},\dots,r_{0}^{M}}$ from $p_{data}\sim\mathcal{D}$.
5      Sample $K$ states for every sample of the minibatch $\{S_{1},\dots,S_{K}\}_{1},\dots,\{S_{1},\dots,S_{K}\}_{M}$} from $q_{\lambda,\hat{\theta}}\propto exp(-U_{\lambda,\hat{\theta}})$ (Using Molecular Dynamics trajectories of length $N$)
@@ 6      Update the neural network parameters by stochastic gradient descent:
$\langle RMSD\rangle\simeq\sum_{i=1}^{K}w_{i}RMSD(\textbf{r}_{0},\textbf{r}_{i})$
$\mathcal{L}_{\theta}=\frac{1}{M}\sum^{M}_{k=1}[max(0,\langle RMSD\rangle+m)]$
@@ 2.3 Markov State Models
Markov State Models (MSM) 48, 49, 50, 51, 52 are employed to analyze the CG simulations and compare them with their corresponding all-atom simulations. MSMs can describe the entire dynamics of a system by partitioning it into $n$ discrete states. For a system to be Markovian, it must be ”memoryless”, meaning that future states depend only on the current state. In the case of Markovian systems, such as MD simulations, a transition probability matrix can be constructed, characterized by the $n$ states and the lag time $\tau$ at which the system’s state is recorded. From this matrix, state populations and conditional pairwise transition probabilities can be obtained, and the state populations can be converted into free energies.
In this work, we employ time-lagged independent component analysis (TICA) 53, 54 to project the high-dimensional conformational space into an optimally reduced low-dimensional space. Following this, the resulting space is discretized using K-means clustering for MSM construction. We featurize the simulation data into pairwise $C_{\alpha}$ distances and use TICA to project the data onto the first four components. For the coarse-grained (CG) data, we adopt the approach presented by Maciej et al. 39, using the covariance matrices of the all-atom molecular dynamics (MD) to project the first three components. This method ensures consistency with established methodologies and facilitates further analysis.
We use MSMs in the coarse-grained trajectories because we had them already for the all-atom trajectories and to evaluate the shape of the folding basins. However, we have no expectation that the thermodynamics or kinetics of these coarse-grained simulations have anything to do with the original ones as the training methods do not preserve these quantities. It can be interpreted as a way to get stable states, which we can take as predictions for the native structure. By comparing the most probable macrostate to the protein’s native conformation, we can evaluate the predictive capabilities of our CG model. For clustering the data we apply the Pairwise Constrained Component Analysis 55 (PCCA) algorithm.
@@ 2.4 Datasets
The first dataset comprises the crystal structures of 12 fast-folding proteins, previously studied by Lindorff-Larsen et al.56 (using all-atom MD) and Majewski et al 39 (using Machine Learned CG MD). These proteins exhibit a variety of secondary structure elements, including $\alpha$-helices and $\beta$-sheets.
The second dataset, used for training the general model, was created by searching the AlphaFold Database 57. This dataset contains 14,871 proteins between 50 and 150 residues, with computationally predicted structures solved using AlphaFold2 58. The selected structures in this dataset are high-quality predictions, with a global predicted local-distance difference test (pLDDT) higher than 90, and are clustered at 50% sequence identity using Usearch 59. We also removed any hits with more than 20% sequence identity to the fast-folding proteins used for testing. This approach ensures a diverse representation of protein structures in the dataset, allowing the NNP to generalize effectively to new and previously unseen protein structures. By incorporating AlphaFold2-predicted structures into our dataset, we significantly increase its size compared to using experimentally solved structures alone.
@@ 3 Results
We trained the FF-NNP using the dataset of 12 fast-folding proteins. For training, the learning rate was set to $\epsilon=0.0001$, and the loss function was defined by Equation (5). The simulation temperature was set to 298 K, and the timestep was 5 fs. We employed trajectories of 1024 steps, with 128 states used for reweighting each trajectory. The margin was set to $m=-1.0;\text{\AA}$, and the mini-batch size was 12. Compared to bottom-up approaches 39, 41, 42, 32, 37, our training process did not require generating expensive reference all-atom data beforehand, and the training took only 5 hours on a single NVIDIA GeForce RTX 2080 GPU. Despite this, the model successfully folded most of the proteins and stabilized their native conformations.
@@
FF-NNP G-NNP All-atom
Protein Min RMSD Mean RMSD Macro prob. Min RMSD Mean RMSD Macro prob. Min RMSD Mean RMSD Macro prob.
(Å) (Å) (%) (Å) (Å) (%) (Å) (Å) (%)
Chignolin 0.3 0.9 $\pm$ 0.7 42.5 $\pm$ 0.2 0.3 1.8 $\pm$ 0.4 24.5 $\pm$ 0.1 0.1 1.0 $\pm$ 0.4 57.5 $\pm$ 0.2
Trp-cage 1.1 3.2 $\pm$ 1.4 5.5 $\pm$ 0.1 1,9 5.2 $\pm$ 1.3 1.7 $\pm$ 0.1 0.4 2.5 $\pm$ 0.8 30.1 $\pm$ 0.2
BBA 0.4 1.3 $\pm$ 0.7 30.0 $\pm$ 0.1 1.7 3.0 $\pm$ 0.5 17.3 $\pm$ 0.1 1.1 3.9 $\pm$ 1.3 5.2 $\pm$ 0.2
WW-domain 0.4 1.0 $\pm$ 0.6 1.8 $\pm$ 0.2 2.5 6.5 $\pm$ 1.1 16.5 $\pm$ 0.1 0.7 2.7 $\pm$ 1.1 45.5 $\pm$ 0.0
Villin 0.4 1.0 $\pm$ 0.5 18.6 $\pm$ 0.1 4.3 7.1 $\pm$ 0.8 30.6 $\pm$ 0.1 0.5 3.4 $\pm$ 1.8 69.4 $\pm$ 0.2
NTL9 0.5 0.9 $\pm$ 0.3 16.0 $\pm$ 0.1 1.6 4.2 $\pm$ 0.8 11.2 $\pm$ 0.1 0.3 1.6 $\pm$ 0.9 15.3 $\pm$ 0.2
BBL 0.5 1.8 $\pm$ 0.6 4.6 $\pm$ 0.2 2.5 6.3 $\pm$ 1.3 36.7 $\pm$ 0.1 1.2 3.1 $\pm$ 0.8 5.2 $\pm$ 0.1
Protein B 4.0 8.6 $\pm$ 2.5 9.9 $\pm$ 0.1 3.7 4.5 $\pm$ 0.5 11.5 $\pm$ 0.2 1.2 4.4 $\pm$ 1.4 30.1 $\pm$ 0.1
Homeodomain 0.5 3.6 $\pm$ 3.9 35.0 $\pm$ 0.1 1.6 6.2 $\pm$ 2.1 39.2 $\pm$ 0.1 0.3 2.3 $\pm$ 1.5 53.5 $\pm$ 0.2
Protein G 0.5 1.0 $\pm$ 0.4 1.5 $\pm$ 0.1 1.3 3.6 $\pm$ 1.0 12.6 $\pm$ 0.1 0.6 2.9 $\pm$ 1.9 17.1 $\pm$ 0.1
a3d 3.4 8.5 $\pm$ 2.3 5.8 $\pm$ 0.1 2.0 3.7 $\pm$ 1.0 29.8 $\pm$ 0.1 1.8 3.5 $\pm$ 0.7 67.9 $\pm$ 0.3
$\lambda$-repressor 4.9 6.8 $\pm$ 1.5 0.4 $\pm$ 0.2 3.6 5.3 $\pm$ 0.6 2.0 $\pm$ 0.2 0.8 4.5 $\pm$ 1.2 21.9 $\pm$ 0.3
Table 1: Minimum average RMSD macrostate statistic from MSM built with CG simulations of the fast-folders, general NNP and all-atom MD. The data shows the average and minimum RMSD of the macrostate, as well as its equilibrium probabilities in percentage (Macro prob.) with standard deviation. In bold the proteins with Mean RMSD $<$ 3.0 Å.
@@
In the previous section, we demonstrated that a folding NNP can be learned for a small set of proteins. In this section, we present the results of the G-NNP, which is trained on a much larger dataset of around 15 thousand protein structures. This enables us to test the generalization capabilities of our approach. We used a random 90/10% training/validation split. Furthermore, we built the training dataset such that it did not contain sequences with a sequence similarity greater than 20% to the fast folders. Thus, we use the 12 fast-folding proteins, providing a well-known but independent test set of our model’s ability to generalize to proteins that fall outside of the training set, allowing us to directly compare the results with the FF-NNP and reference all-atom simulations.
@@
Results from the MSM analysis reveal that for specific proteins, such as Chignolin or BBA, the minimum RMSD macrostate aligns with the native structure, exhibiting an average RMSD of less than 3.0 Å. Moreover, their equilibrium probabilities are 24.5 $\pm$ 0.1 % for Chignolin and 17.3 $\pm$ 0.1 % for BBA (Table 1). This evidence suggests that our NNP can effectively generalize when trained on an ample dataset. For the other proteins, the minimum RMSD of the macrostates is consistently lower than 5 Å. As a result, although the macrostates do not precisely match the native structure, near-native conformations are sampled during the simulations with notable probability.
Simulations for Chignolin, BBA, and Homeodomain proteins successfully sampled folding events, even when starting from completely unfolded conformations. The folding events observed in our simulations, illustrated in Figure 2, provide compelling evidence that our G-NNP generalizes beyond the training set, accurately folding proteins not included in the training set, albeit with limitations in accurately reproducing the thermodynamic landscapes. As illustrated in Figure S3 of the supplementary information, the free-energy landscapes produced by our models do not reproduce those generated by all-atom MD.